In Desmos, type in both equations:

  1. y= square root of x
  2. y= square root x-3+4

What transformations occurred in the graph?


1
Expert's answer
2021-03-17T07:07:11-0400
"y=5\sqrt[3]{x+1}"

Parent function "f(x)=\sqrt[3]{x}."

To graph "g(x)=\sqrt[3]{x+1}," we shift the graph of "f" to the left 1 unit.

To graph "h(x)=5\sqrt[3]{x+1}," stretch the graph of "g" vertically by a factor of "5."Function is given as:

"y=\sqrt[2]{x^{-2}+4}"


"y=\sqrt[2]{\frac{1}{x^2}+4}"


for real solutions of the term inside the square root should be as:

"\frac{1}{x^2}+4\ge0"


We can se that has degree two so for all values of x the above term gives always positive values except x=0, as this makes it infinite value i.e. undefined, so the domain of the given equation is as:

Domain = ("-\infty,+\infty" )-[0] i.e the transformation will lie on x axis